Jake spent 3/8 of his money on videotape and 1/2 of what was left on a cassette Tape. he now has $6.25. How much money did he start with?
Answer by mananth(16946)

Let jake have $x
3x/8 on video tape
Balance left = x-3x/8
=5x/8
He spent (1/2) *(5x/8) on cassette tape
=(5x/16)
So he had a balance of (5x/8)-(5x/16)
=5x/16
6.25 = 6 1/4 = 25/4
5x/16 = 25/4
x= (16*25)/(4*5)
x= 20
He had $20